November 1996  ICTCM IX (Reno)  2-1

Introduction

• instructor must stay within comfort zone (may need to actively work to extend comfort zone, but should not go beyond it when teaching)

• this is easier than it might seem; many good problems can be found in the textbooks

• this talk will provide an example of how this might look in a sophomore-level ODE course